I am having trouble with this problem. I have tried to solve it and I must be missing something in the equation. I have 2X+X-20+25=180. I know this is wrong .
A stairway railing has a design that contains triangles. In one of the triangles one angle measure 20degrees less than twice the smallest angle. The third angle measures 25degrees greater than twice the smallest angle. Determine the measures of the 3 angles.

Answer by stanbon(75887)

Let the smallest angle have measure "x"
Then a 2nd angle is 2x-20
And a 3rd angle is 2x+25
EQUATION:
angle + angle + angle = 180
x+2x-20+2x-25=180
5x+5=180
5x=175
x=35 (measure of smallest angle)
2x-20=50 (measure of 2nd angle)
2x+25=95 (measure of 3rd angle)
Cheers,
Stan H.